If $4,500.00 is investment at 4% simple interest, how much needs to be invested at 10% simple interest so the total interest per year is $730.00.
You need to invest ? at 10% simple interest.

Answer by mananth(16949)

$4500 at4%
interest = 0.04*4500
Interest = $180
Total interest required = $ 730
730-180= $550
550 = x*0.1*1
550/0.1=x
= $5500 @ 10 %
CHECK
5500*0.1+4500*0.04 = $730